Circular No. 8196

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hu Further to IAUC 8191, M. Armstrong reports the discovery of an ap-

parent supernova (mag 18.0) on unfiltered CCD images taken on Sept. 6.968 and 7.919 UT with a 0.35-m reflector. The new object is located at α = 19^{h}11^{m}31^{s}.40, δ = +77^{o}53′35′′.2 (equinox 2000.0), which is ≈ 4′′ west and 2′′.5 south of the center of a galaxy that itself is 204′′ west and 120′′ south of UGC 11423. Nothing was visible at the position of SN 2003hu on Armstrong’s images taken on Aug. 4 and 31 (limiting mag 19.5) or on Palomar Sky Survey red and blue plates.

SUPERNOVAE 2003he AND 2003hs R. J. Foley, R. Chornock, and A. V. Filippenko, University of Califor-

nia, Berkeley; and F. J. D. Serduke, Lawrence Livermore National Laboratory, report that inspection of CCD spectra (range 310–1000 nm), obtained on Sept. 3 UT with the Shane 3-m telescope at Lick Observatory, reveal that SN 2003he (IAUC 8182) is of type Ia, with a spectral-feature age of 9 ± 2 days past maximum brightness; the continuum is quite blue, contrary to the description given on IAUC 8189 (where galaxy contamination was not properly taken into account). SN 2003hs (IAUC 8191) is also of type Ia, but near maximum light. The spectrum exhibits relatively strong 580 nm Si II/Ti II absorption relative to the 615-nm Si II absorption, suggesting that SN 2003hs may be slightly subluminous.

COMET C/2002 T7 (LINEAR)

Ephemeris (H = 5.5, n = 3) from orbital elements on MPEC 2003-R06:

2003 TT α_{2000} δ_{2000} Δ r ϵ β Mag.
Aug. 29 5^{h}38.^{m}32 +31^{o}12.^{′}0 3.931 3.711 70.^{o}1 14.^{o}8 12.7
Sept. 8 5 40.10 +31 49.0 3.643 3.595  79.3 16.0 12.5
18 5 39.95 +32 32.2 3.348 3.477  88.9 16.8 12.2
28 5 37.22 +33 22.6 3.050 3.358  99.1 17.1 11.9
Oct. 8 5 30.97 +34 20.7 2.756 3.237 110.0 16.9 11.5
18 5 20.05 +35 25.3 2.474 3.114 121.6 15.8 11.2
28 5 02.96 +36 31.3 2.211 2.990 134.2 13.8 10.8
Nov. 7 4 38.14 +37 26.5 1.978 2.863 147.4 10.7 10.4
17 4 04.77 +37 48.7 1.787 2.735 159.3  7.3 10.0
27 3 24.07 +37 08.7 1.650 2.604 161.1 7.0 9.7
Dec. 7 2 40.40 +35 06.6 1.576 2.471 148.7 12.0  9.4
2003 September 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green